Can you help me, please??? this is our assignment and I can't figure it out..
The question is :
Henry's father is now twice as old as Henry. Sixteen years ago, he was four times as old. How old is each now?

Thank you. .

Answer by stanbon(29602)

Henry's father is now twice as old as Henry. Sixteen years ago, he was four times as old. How old is each now?
------------
NOW DATA:
Let Henry's age be "x" ; Father is "2x".
--------------
16 yrs. ago DATA:
Henry was x-16 ; Father was 2x-16
-----------------
EQUATION:
Father was four times as old.
2x-16 = 4(x-16)
2x-16 = 4x-64
2x = 48
x=24 (Henry's age NOW)
2x=48 (Father's age NOW)
====================
Cheers,
Stan H.